OpenStudy (chillout):

Solve the initial value problem: ty' + 2y = sin(t) -> y(π/2)=1, t>0

OpenStudy (chillout):

I've just started studying ODEs, so it'd great if anyone solved it by integrating factors method

OpenStudy (anonymous):

\[\Large y'+\frac{2}{t} y=\frac{\sin(t)}{t}\] Can you tell from this how an integration factor must look like?

OpenStudy (chillout):

If I had to guess (I'm completely new to this), I'd guess 2/t.

OpenStudy (anonymous):

Well, basically there is (as for many things) a formula to find this integrating factor: \[\Large \mu(x)=\exp\left(\int p(x)dx \right)\] Where \(p(x)\) is the function in front the linear term when the DE is in standard form: \[\Large y'+p(x)y=q(x) \]

OpenStudy (anonymous):

most people compute it via the formula, because it's easy to remember, if you want to derive it yourself though by doing it manually, we could take a look at this too.

OpenStudy (anonymous):

The basic idea behind is that the LHS of the DE looks almost like a product rule of differentiation, integrating factors take advantage of that.

OpenStudy (chillout):

I noticed it. Does this mean that I'll have to do this? \[μ(x)*\frac{ dy }{ dx } + p(x)*μ(x)*y(x) = q(x)*μ(x)\]

OpenStudy (anonymous):

exactly, do you want to do the general case or applied to your problem already? But you're on the right track anyway, the next thing you would want to do is compute \[\Large\left(\mu(x)y\right)'=\mu(x)'y+y'\mu(x) \] And then match coefficients

OpenStudy (chillout):

With the general case explanation I think I got it.

OpenStudy (chillout):

Just one moment 'till I solve it.

OpenStudy (anonymous):

very good then :-) ! no hurries.

OpenStudy (chillout):

All right, Now I have that (yt²)' = t*sin(t). I suppose now I'll just have to integrate both sides?

OpenStudy (anonymous):

exactly, you got it!

OpenStudy (chillout):

The solution for this DE is \[\frac{ \sin(t) - t*\cos(t) + C }{ t² }\] putting in the value for y gives C = 1/4(pi²-4)

OpenStudy (anonymous):

looks right to me

OpenStudy (chillout):

Reached the answer, many thanks!

OpenStudy (anonymous):

you are welcome